# Ramulator 2.0: A Modern, Modular, and Extensible DRAM Simulator

### v1

@@ 1 Introduction
In summary, the key features and contributions of Ramulator 2.0 are:
  • Ramulator 2.0 is a modular and extensible DRAM simulator written in C++20 [11] that enables rapid and agile implementation and evaluation of design changes in the memory system. Ramulator 2.0 can either work as a standalone simulator, or be used as a memory system library by a system simulator (e.g., gem5 [18], ZSim [19]).
@@ 2.1 Modular and Extensible Software Architecture
Figure 1 shows the high-level software architecture of Ramulator 2.0 with the key interfaces we identify in a DRAM-based memory system (dark boxes) and their typical implementations (light boxes) when modeling a DDR5 system with RowHammer mitigation. The arrows illustrate the relationships among different components in the simulated system (i.e., how they call each other’s interface functions). We highlight the memory request path with red arrows, DRAM command path with blue arrows, and DRAM maintenance requests (e.g., refreshes) with green arrows. A typical execution of the simulation is as follows: First, memory requests are sent 1 from the frontend (either parsed from traces or generated by another simulator, e.g., gem5 [18]) to the memory system, where the memory addresses are mapped 2 to the DRAM organization through the address mapper. Then, the requests are enqueued 3 in the request buffers of the DRAM controller. The DRAM controller is responsible for 1) ticking the refresh manager 4, which could enqueue high-priority maintenance requests (e.g., refreshes) back to the controller, 2) querying the request scheduler 5, which in turn consults the DRAM device model 6 to decode the best DRAM command to issue 7 to serve a memory request, and 3) issuing the DRAM command 8, which updates the behavior and timing information of the DRAM device model. Finally, the memory controller executes the finished request’s callback 9 to notify the frontend.
@@ 2.1.1 Memory Controller Plugins
We make a key observation that many modeled functions in the memory controller (e.g., controller-based RowHammer mitigations that tracks the issued activation commands) and utilities needed for evaluation (e.g., collecting statistics from the issued DRAM commands and analyzing the memory access patterns) are triggered (updated) by the currently-scheduled DRAM command. To avoid having many similar memory controller implementations for every single such modeled function and utility, we model these functions as plugins to the memory controller. As an example, Figure 2 shows in detail how various RowHammer mitigation techniques (e.g., PARA [12], Graphene [14], Hydra [15]) can be implemented as such controller plugins.
@@
The plugin interface has a simple update(DRAM_CMD, ADDR) function that the controller calls ( 1 in Figure 1 and 2) to notify the plugin implementations about the DRAM command and address issued by the memory controller. The RowHammer mitigation implementation then updates its internal state (e.g., generates a random number for PARA, updates the row activation count table for Graphene, or queries the row count cache for Hydra). If the implementation detects the need to refresh the potential RowHammer victim rows, it calls the priority_enqueue() function ( 2 in Figure 1 and 2) of the memory controller interface to send a high-priority refresh request for the identified victim rows, ready to be scheduled in the following cycles, as determined by the mitigation techniques. To showcase the modularity and extensibility of memory controller plugins, Section 3.3 provides a cross-sectional evaluation of the performance overhead of six different RowHammer mitigation techniques, all implemented as memory controller plugins.
@@ 2.2 Concise and Intuitive DRAM Specifications
Second, Ramulator 2.0 implements the DRAM commands (e.g., the state changes caused by the DRAM commands and the prerequisite commands based on the current state) using a library of lambda functions. These functions are implemented in a templated way so that they are defined only once, but can be reused many times for similar DRAM commands across different standards. As an example, Listing 3 shows a part of implementations of the RFMab command (all-bank refresh management, which exists in the DDR5 [7], LPDDR5 [8], GDDR6 [9], and HBM3 [10] standards) that requires all the banks in the DRAM device to be closed before it can be issued.
@@
In the lambda function library source file (i.e., lambdas/preq.h), Ramulator 2.0 defines a generic function RequireAllBanksClosed(...) that checks for all banks in the organization hierarchy if all of them are closed (lines 6-7). If so, it simply returns the input command argument cmd (line 12), indicating that no prerequisite command is needed for cmd. Otherwise, it returns the PREab (precharge all-bank) command to close all the banks first. This function is templated on the DRAM standard implementation (i.e., the DRAM_t template parameter on line 2) so that it can automatically get the correct integer encoding of the commands and states for different DRAM standards at compile-time. By reusing this templated function in different DRAM standards (lines 16, 18, 20, and 22), implementing the prerequisite checks for the RFMab command needs only a single line of code in each standard (instead of duplicating the entire RequireAllBanksClosed function for each DRAM standard as in Ramulator 1.0).
@@ 3 Validation & Evaluation
We compare the simulation speed of Ramulator 2.0 with three other cycle-accurate DRAM simulators: Ramulator 1.0 [2], DRAMsim2 [3], DRAMsim3 [4], and USIMM [1]. All four simulators are configured with comparable system parameters. We generate two memory traces, one with a random access pattern and another with a streaming access pattern, each containing five million memory requests (read-write ratio = 4:1). For each simulator and trace, we repeat the simulation for each trace ten times. Table I shows the minimum, average, and maximum simulation runtimes across the ten repetitions. We conclude that, despite the increased modularity and extensibility, Ramulator 2.0 achieves a comparably fast (and even faster) simulation speed versus other existing cycle-accurate DRAM simulators. We provide the complete set of scripts, configurations, and traces to reproduce our results in [6].
@@ 3.3 Cross-Sectional Study of RowHammer Mitigations
To demonstrate the modularity and extensibility of Ramulator 2.0, we implement six different RowHammer mitigation techniques, PARA [12], an idealized version of TWiCe [13], Graphene [14], Hydra [15], Randomized Row-Swap (RRS) [16], and an ideal refresh-based mitigation (Ideal) [17]. All of these mechanisms are implemented in the form of memory controller plugins as described in Section 2.1.1. Figure 3 shows the performance overhead (weighted speedup normalized to a baseline configuration running the same workloads without any RowHammer mitigation, y-axis) of different RowHammer mitigations as the RowHammer threshold (i.e., the minimum number of DRAM row activations to cause at least one bitflip, tRH, x-axis) decreases from 5000 to 10. We use traces generated from SPEC2006 [23] and SPEC2017 [24] to form 25 four-core multiprogrammed workloads that we feed through a simplistic out-of-order core model (the complete set of scripts and traces to reproduce these experiments are in [6]).
We make the following two observations. First, all evaluated RowHammer mitigations (except for Ideal) cause significant performance overhead compared to the ideal mitigation as tRH decreases to very low values. Second, for tRH <50<50, the performance overhead of RRS becomes too high for the simulation to make progress. The reason for this is that the activation caused by a row swap triggers even more row swaps, preventing DRAM from serving memory access requests. We conclude that existing RowHammer mitigation techniques are not scalable enough to very low tRH values (below 50). As such, more research effort is needed to develop more efficient and scalable RowHammer mitigation techniques.

### v2

@@ 1 Introduction
The key features and contributions of Ramulator 2.0 are:
  • Ramulator 2.0 is a modular and extensible DRAM simulator written in C++20 that enables rapid and agile implementation and evaluation of design changes in the memory system. Ramulator 2.0 can either work as a standalone simulator, or be used as a memory system library by a system simulator (e.g., gem5 [18], zsim [19]).
@@ 2.1 Modular and Extensible Software Architecture
Figure 1 shows the high-level software architecture of Ramulator 2.0 with the key interfaces we identify in a DRAM-based memory system (dark boxes) and their typical implementations (light boxes) when modeling a DDR5 system with RowHammer mitigation. The arrows illustrate the relationships among different components in the simulated system (i.e., how they call each other’s interface functions). We highlight the memory request path with red arrows, DRAM command path with blue arrows, and DRAM maintenance requests (e.g., refreshes) with green arrows. A typical execution of the simulation is as follows: First, memory requests are sent 1 from the frontend (either parsed from traces or generated by another simulator, e.g., gem5 [18]) to the memory system, where the memory addresses are mapped 2 to the DRAM organization through the address mapper. Then, the requests are enqueued 3 in the request buffers of the DRAM controller. The DRAM controller is responsible for 1) ticking the refresh manager 4, which could enqueue high-priority maintenance requests (e.g., refreshes) back to the controller, 2) querying the request scheduler 5, which in turn consults the DRAM device model 6 to decode the best DRAM command to issue 7 to serve a memory request, and 3) issuing the DRAM command 8, which updates the behavior and timing information of the DRAM device model. Finally, the memory controller executes the finished request’s callback 9 to notify the frontend of the completion of the memory request.
@@ 2.1.1 Memory Controller Plugins
We make a key observation that many modeled functions in the memory controller (e.g., controller-based RowHammer mitigations that tracks the issued activation commands) and utilities needed for evaluation (e.g., collecting statistics from the issued DRAM commands and analyzing the memory access patterns) are triggered (updated) by the currently-scheduled DRAM command. To avoid having many similar memory controller implementations for every single such modeled function and utility, we model these functions as plugins to the memory controller. As an example, Figure 2 shows in detail how various RowHammer mitigation techniques (e.g., PARA [12], Graphene [14], Hydra [15], TRR [22, 23], RFM [7]) can be implemented as such memory controller plugins.
@@
The plugin interface has a simple update(DRAM_CMD, ADDR) function that the controller calls ( 1 in Figure 1 and 2) to notify the plugin implementations about the DRAM command and address issued by the memory controller. The RowHammer mitigation implementation then updates its internal state (e.g., generates a random number for PARA, updates the row activation count table (bank activation counter) for Graphene and TRR (RFM), or queries the row count cache for Hydra). If the implementation detects the need to refresh the potential RowHammer victim rows, it calls the priority_enqueue() function ( 2 in Figure 1 and 2) of the memory controller interface to send a high-priority refresh request for the identified victim rows, ready to be scheduled in the following cycles, as determined by the mitigation techniques. To showcase the modularity and extensibility of memory controller plugins, Section 3.3 provides a cross-sectional evaluation of the performance overhead of six different RowHammer mitigation techniques, all implemented as memory controller plugins.
@@ 2.2 Concise and Intuitive DRAM Specifications
Second, Ramulator 2.0 implements the DRAM commands (e.g., the state changes caused by the DRAM commands and the prerequisite commands based on the current state) using a library of lambda functions. These functions are implemented in a templated way so that they are defined only once, but can be reused many times for similar DRAM commands across different standards. As an example, Listing 3 shows a part of implementations of the RFMab command (all-bank refresh management, which exists in the DDR5 [7], LPDDR5 [8], GDDR6 [9], and HBM3 [10] standards) that requires all the banks to be closed before it can be issued.
@@
Ramulator 2.0 defines a RequireAllBanksClosed generic function that checks for all banks in the organization hierarchy if all of them are closed (lines 6-7). If so, it simply returns the input command argument cmd (line 12), indicating that no prerequisite command is needed for cmd. Otherwise, it returns the PREab (precharge all-bank) command to close all the banks first. This function is templated on the DRAM standard implementation (i.e., the DRAM_t template parameter on line 2) so that it can automatically get the correct integer encoding of the commands and states for different DRAM standards at compile-time. By reusing this templated function in different DRAM standards (lines 16, 18, 20, and 22), implementing the prerequisite checks for the RFMab command needs only a single line of code in each standard (instead of duplicating the entire RequireAllBanksClosed function for each DRAM standard as in Ramulator 1.0).
@@ 3 Validation & Evaluation
We compare the simulation speed of Ramulator 2.0 with three other cycle-accurate DRAM simulators: Ramulator 1.0 [2], DRAMsim2 [3], DRAMsim3 [4], and USIMM [1]. All four simulators are compiled with gcc-12 -O3, and configured with comparable system parameters. We generate two memory traces, one with a random access pattern and another with a streaming access pattern, each containing five million memory requests (read-write ratio = 4:1). For each simulator and trace, we run the simulation for each trace ten times on a machine with an Intel Xeon Gold 5118 processor. Table I shows the minimum, average, and maximum simulation runtimes across the ten runs. We conclude that, despite the increased modularity and extensibility, Ramulator 2.0 achieves a comparably fast (and even faster) simulation speed versus other existing cycle-accurate DRAM simulators. We provide the scripts, configurations, and traces to reproduce our results in Ramulator 2.0’s repository [6].
@@ 3.3 Cross-Sectional Study of RowHammer Mitigations
To demonstrate the modularity and extensibility of Ramulator 2.0, we implement six different RowHammer mitigation techniques, PARA [12], an idealized version of TWiCe [13], Graphene [14], Hydra [15], Randomized Row-Swap (RRS) [16], and an ideal refresh-based mitigation (Ideal) [17]. All of these mechanisms are implemented in the form of memory controller plugins as described in Section 2.1.1. Figure 3 shows the performance overhead (weighted speedup normalized to a baseline configuration running the same workloads without any RowHammer mitigation, y-axis) of different RowHammer mitigations as the RowHammer threshold (i.e., the minimum number of DRAM row activations to cause at least one bitflip, $\mathrm{t}_{\mathrm{RH}}$, x-axis) decreases from 5000 to 10. We use traces generated from SPEC2006 [25] and SPEC2017 [26] to form 25 four-core multiprogrammed workloads that we feed through a simplistic out-of-order core model (the complete set of scripts and traces to reproduce these experiments are in [6]).
We make the following two observations. First, all evaluated RowHammer mitigations (except for Ideal) cause significant performance overhead compared to the ideal mitigation as $\mathrm{t}_{\mathrm{RH}}$ decreases to very low values. Second, for $\mathrm{t}_{\mathrm{RH}}$ <50<50, the performance overhead of RRS becomes too high for the simulation to make progress. This is because the activation caused by a row swap triggers even more row swaps, preventing DRAM from serving memory access requests. We conclude that existing RowHammer mitigation techniques are not scalable enough to low $\mathrm{t}_{\mathrm{RH}}$ values (<<50). As such, more research effort is needed to develop more efficient and scalable RowHammer mitigation techniques.